The school flagpole cast a shadow 15ft long. Mr. Hansen cast a shadow 4ft long. Mr. Hansen is 6ft tall. How tall is the flag pole.
Answer by checkley75(3666)

15/6=X/4 CROSS MULTIPLY.
6X=60
X=60/6
X=10 FEET FOR THE FLAGPOLE SHADOW.